\begin{thebibliography}{AAAA}

\bibitem[1]{BL} F.A.Berezin, D.A.Leites, {\em Supermanifolds}, Soviet
Math. Dokl. \textbf{16} (1975), 1218-1222.

\bibitem [2] {BK}
F.A. Berezin, G.I. Kac, {\em
Lie groups with commuting and anticommuting parameters.} (Russian)
Mat. Sb. (N.S.) \textbf{82} (124) 1970 343--359. 

\bibitem [3] {Be} Berezin, F. A. {\em Representations of the supergroup $U(p,q).$} (Russian)  Funkcional. Anal. i Prilo\v zen.  10  (1976), no. \textbf{3}, 70--71.


\bibitem[4]{Be-Re} F.A. Beresin, V.S. Retakh, {\em A method of computing characteristic classes of vector bundles.}  Rep. Math. Phys.  18  (1980),  no. \textbf{3}, 363--378. 

\bibitem [5] {Be2} F.A. Berezin, {\em Introduction to superanalysis.} Edited and with a foreword by A. A. Kirillov. With an appendix by V. I. Ogievetsky. Translated from the Russian by J. Niederle and R. Koteck\'y. Translation edited by Dimitri Le\u\i tes. Mathematical Physics and Applied Mathematics, \textbf{9}. D. Reidel Publishing Co., Dordrecht, 1987

\bibitem[6]{Ch-Si} J. Cheeger, J. Simons, {\em
Differential characters and geometric invariants}, Geometry and topology (College Park, Md., 1983/84), 
50--80, Lecture Notes in Math., \textbf{1167}, Springer, Berlin, 1985. 

\bibitem[7]{Chn-Si} S.S. Chern, J. Simons, {\em Characteristic forms and geometric invariants},  Ann. of Math. (2)  \textbf{99}  (1974), 48--69. 


\bibitem[8]{De-Mo} P. Deligne, J.H. Morgan, {\em Notes on supersymmetry }(following Joseph Bernstein).  Quantum fields and strings: a course for mathematicians, Vol. \textbf{1}, \textbf{2} (Princeton, NJ, 1996/1997),  41--97, Amer. Math. Soc., Providence, RI, 1999.

\bibitem[9]{DHZ} J. Dupont, R. Hain, S. Zucker, {\em Regulators and characteristic classes of flat bundles},  The arithmetic and geometry of algebraic cycles (Banff, AB, 1998),  47--92, CRM Proc. Lecture Notes, \textbf{24}, Amer. Math. Soc., Providence, RI, 2000. 

\bibitem[10]{Esnault} H. Esnault, {\em Characteristic classes of flat bundles}, Topology  27  (1988),  no. \textbf{3}, 323--352.

\bibitem[11]{Iy-Si} J.N. Iyer, C.T. Simpson, {\em Regulators of canonical extensions are torsion; the smooth divisor case}, preprint 2007, arXiv math.AG/07070372.

\bibitem[12]{Ka} V.G. Kac, {\em Characters of typical representations of classical Lie superalgebras.}  Comm. Algebra  5  (1977), no. \textbf{8}, 889--897.

\bibitem[13]{Ka2} V.G. Kac, {\em Lie superalgebras.}  Advances in Math.  26  (1977), no. \textbf{1}, 8--96.

\bibitem[14]{La} C. Laurent-Gengoux, {\em Secondary characteristic classes of super-foliations.}  C. R. Math. Acad. Sci. Paris  339  (2004),  no. \textbf{8}, 567--572.

\bibitem[15]{Le} D.A. Leites, {\em  Introduction to the theory of supermanifolds.} (Russian)  Uspekhi Mat. Nauk  35  (1980), no. \textbf{1} (211), 3--57, 255.
We used the yet unpublished book {\em Seminar on Supersymmetries} (English)
by the same author.


\bibitem[16]{Ma} Y.I. Manin, {\em Gauge field theory and complex geometry.} Translated from the Russian by N. Koblitz and J. R. King. Grundlehren der Mathematischen Wissenschaften [Fundamental Principles of Mathematical Sciences], \textbf{289}. Springer-Verlag, Berlin, 1988. x+297 pp.


\bibitem[17]{Narasimhan} M. S. Narasimhan, S. Ramanan, {\em Existence of universal connections},  Amer. J. Math.  \textbf{83}  1961 563--572. 

\bibitem[18]{Narasimhan2} M. S. Narasimhan, S. Ramanan, {\em Existence of universal connections. II},  Amer. J. Math.  \textbf{85}  1963 223--231.


\bibitem[19]{Quillen} D. Quillen, {\em Superconnections and the Chern character.}  Topology  24  (1985),  no. \textbf{1}, 89--95.

 \bibitem[20]{Quillen2} D. Quillen, {\em Superconnection character forms and the Cayley transform.}  Topology  27  (1988),  no. \textbf{2}, 211--238. 

\bibitem [21]{Reznikov} A. Reznikov, {\em Rationality of Chern-Simons classes},  J. Differential Geom.  43  (1996),  no. \textbf{3}, 674--692. 


\bibitem [22]{Reznikov2} A. Reznikov,  {\em All regulators of flat bundles are torsion}, Ann. of Math. (2) 141 (1995), no. \textbf{2}, 373--386.


\bibitem [23]{Se} A. Sergeev, {\em The invariant polynomials on simple Lie superalgebras.}  Represent. Theory  \textbf{3}  (1999), 250--280 

\bibitem [24]{Sh} V. Shander {\em Invariant functions on supermatrices}, arXiv:math/9810112. 


\bibitem[25]{Sm-Su} J. Simons, D. Sullivan, {\em Axiomatic Characterization of Ordinary Differential Cohomology}, arXiv math.AT/0701077.


\end{thebibliography}

\documentclass[12pt]{amsart}

\newtheorem{theorem}{Theorem}[section]
\newtheorem{lemma}[theorem]{Lemma}

\newtheorem{proposition}[theorem]{Proposition}
\newtheorem{remark}[theorem]{Remark}

\newcommand{\ncom}{\newcommand}
\ncom{\ep}{\epsilon}
\ncom{\rar}{\rightarrow}
\ncom{\lrar}{\longrightarrow}
\ncom{\ov}{\overline}
\ncom{\what}{\widehat}

\newcommand{\ignore}[1]{}
\ncom{\m}{\mbox}
\ncom{\sta}{\stackrel}
\ncom{\comx}{{\mathbb C}}
\ncom{\A}{{\mathbb A}}
\ncom{\Z}{{\mathbb Z}}
\ncom{\Q}{{\mathbb Q}}
\ncom{\R}{{\mathbb R}}
\ncom{\G}{{\mathbb G}}
\ncom{\hH}{{\mathbb H}}
\ncom{\al}{\alpha}
\ncom{\p}{{\mathbb P}}
\ncom{\E}{{\mathbb E}}
\ncom{\N}{{\mathbb N}}
\ncom{\K}{{\mathbb K}}
\ncom{\X}{{\mathbb X}}
\ncom{\f}{\frac}
\ncom{\cA}{{\mathcal A}}
\ncom{\cB}{{\mathcal B}}
\ncom{\cD}{{\mathcal D}}
\ncom{\cX}{{\mathcal X}}
\ncom{\cO}{{\mathcal O}}
\ncom{\cW}{{\mathcal W}}
\ncom{\cL}{{\mathcal L}}
\ncom{\cP}{{\mathcal P}}
\ncom{\cH}{{\mathcal H}}
\ncom{\cS}{{\mathcal S}}
\ncom{\cM}{{\mathcal M}}
\ncom{\cC}{{\mathcal C}}
\ncom{\cT}{{\mathcal T}}
\ncom{\cF}{{\mathcal F}}
\ncom{\cN}{{\mathcal N}}
\ncom{\cJ}{{\mathcal J}}
\ncom{\cV}{{\mathcal V}}
\ncom{\cZ}{{\mathcal Z}}
\ncom{\cU}{{\mathcal U}}
\ncom{\cSU}{{\mathcal S \mathcal U}}
\ncom{\cG}{{\mathcal G}}
\ncom{\cQ}{{\mathcal Q}}
\ncom{\cR}{{\mathcal R}}
\ncom{\cY}{{\mathcal Y}}
\ncom{\cE}{{\mathcal E}}
\ncom{\cI}{{\mathcal I}}
\ncom{\mylabel}[1]{{\rm (#1)}\label{#1}}
\ncom{\Hom}{{\textit{Hom}}}
\ncom{\eop}{{\hfill $\Box$}}
\begin{document}
\baselineskip=16pt

%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%

\title[Chern-Simons classes]{Chern-Simons classes of flat connections on supermanifolds }

%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%

\author[J. N. Iyer]{Jaya NN  Iyer}
\address{School of Mathematics, Institute for Advanced Study, 1 Einstein Drive, Princeton NJ 08540 USA.}
\email{jniyer@ias.edu}

\address{The Institute of Mathematical Sciences, CIT
Campus, Taramani, Chennai 600113, India}
\email{jniyer@imsc.res.in}

\author[U.N.Iyer]{Uma N Iyer}
\address{308A, Department of Mathematics and Computer Science,
CP315, Bronx Community College,
University Avenue and West 181 Street,
Bronx, NY 10453.}
\email{uma.iyer@bcc.cuny.edu}

\footnotetext{Mathematics Classification Number: 53C05, 53C07, 53C29 }
\footnotetext{Keywords: Supermanifolds, Connections, Secondary classes.}

\begin{abstract}
In this note we define Chern-Simons classes of a superconnection $D+L$ on a complex supervector bundle $E$ such that $D$ is flat and preserves the grading, and $L$ is an odd endomorphism of $E$ on a supermanifold. As an application we obtain a definition of Chern-Simons classes of a (not necessarily flat) morphism between flat vector bundles on a smooth manifold. 
We extend Reznikov's theorem on triviality of these classes 
when the manifold is a compact K\"ahler manifold or a smooth complex quasi--projective variety, in degrees $>\,1$.
\end{abstract}
\maketitle

%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%

\section{Introduction}

%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%

Suppose $(X,\cC^{\infty}_X)$ is a $\cC^{\infty}$-differentiable manifold endowed with the structure sheaf
$\cC^{\infty}_X$ of smooth functions.
Let $E$ be a complex $\cC^{\infty}$ vector bundle on $X$ of rank $r$ and equipped with a connection $\nabla$.
The Chern-Weil theory defines the Chern classes
$$
c_i(E,\nabla) \in H^{2i}_{dR}(X,\comx), \m{  for }i=0,1,...,r
$$
in the de Rham cohomology of $X$.
These classes are expressed in terms of the $GL_r$-invariant polynomials evaluated on the curvature form
$\nabla^2$.

Suppose $E$ has a flat connection, i.e., $\nabla^2=0$. Then the de Rham Chern classes are zero. It is significant to define Chern-Simons classes for a flat connection. These are classes in the $\comx/\Z$-cohomology and were defined by Chern-Cheeger-Simons in \cite{Ch-Si}, \cite{Chn-Si}.   

Quillen has pointed out in \cite{Quillen},\cite{Quillen2}, a homomorphism 
$u:E_0\rar E_1$ between vector bundles on a smooth manifold $M$ and inducing an isomorphism over a subset 
$A\subset M$ corresponds to an element in the relative $K$-group $K(M,A)$. A Chern character in the de Rham cohomology of $M$ associated to the homomorphism $u$ is computed in \cite{Quillen} whose class is shown to be equal to the difference $\m{ch}(E_0)- \m{ch}(E_1)$ of the Chern characters. This describes the Chern character of the homomorphism $u$. In fact, we think that it would be good to look at a quiver, i.e., a sequence of homomorphisms between vector bundles 
$$
E_0\rar E_1\rar...\rar E_r
$$
over a smooth manifold and define the Chern character of the sequence in the de Rham cohomology. This will involve a study of $\Z_{r+1}$-graded objects, which we will look in the future. 
Quillens proof involves
regarding $E=E_0\oplus E_1$ as a 
supervector bundle on $M$ and $D$ be any connection preserving the grading and associating an odd endomorphism of $E$, with respect to $u$ and a choice of a metric. 

In this paper, we want to look at a morphism $u$ between flat vector bundles and extend Quillen's
construction and define Chern-Simons classes for the morphism $u$. 
Hence it is relevant to define Chern-Simons classes for flat connections in the setting of supermanifolds, in a more general set-up.

For the definition of \textit{supermanifolds}, due to F.A.Berezin and D. Leites, see \cite{De-Mo} (as well as \cite{BL}, \cite{Le}).
The Chern classes of supervector bundles are defined in \cite{Be-Re}, 
on a supermanifold in the integral cohomology. We note that the usual Chern-Weil theory on smooth manifolds expresses de Rham Chern classes in terms of $GL_n$-invariant polynomials on the curvature form of a connection on a smooth vector bundle. In the supersetting, a study of the invariant polynomials has been carried out
by Sergeev \cite{Se}, following works by Berezin \cite{Be},\cite{Be2} and Kac \cite{Ka}. We do not know if the differential forms defined using the invariant polynomials of Sergeev give the de Rham Chern class of a super vectorbundle, as obtained by Quillen.
In this paper we consider the Chern character defined by Quillen to define the Chern-Simons classes.

Let $(M,\cO_M)$ denote a complex supermanifold and $(M_B,\cC^\infty_M)$ 
denote the underlying $\cC^\infty$-manifold.

With notations as in \cite{De-Mo} or \S \ref{prelim}, we show
\begin{theorem}
Suppose $(M,\cO_M)$ is a complex supermanifold. Let $\cE^{r|s}$ be a complex 
supervector bundle on $(M,\cO_M)$ equipped with a superconnection $\nabla=D+L$ such that $D$ preserves the grading and is flat, and $L$ is an odd endomorphism of $\cE^{r|s}$. 
Then there exists uniquely determined Chern-Simons classes
$$
\widehat{c_i}( \cE^{r|s},\nabla)\, \in \, H^{2i-1}(M_B,\comx/\Z)
$$
for $i>0$. Furthermore, if $M_B$ is a compact K\"ahler manifold or a smooth complex quasi--projective variety, then these classes are 
torsion, in degrees $>\,1$.
\end{theorem}

This can be thought of as an extension of Reznikov's 
fundamental theorem \cite{Reznikov2} on rationality of Chern-Simons classes on compact K\"ahler manifold, in the setting of supermanifolds.   
We also define Chern-Simons classes of a (not necessarily flat) homomorphism $u:E_0\rar E_1$ between flat complex vector bundles, extending Quillen's construction of the de Rham Chern character.
Then we prove a relative Reznikov theorem (see Theorem \ref{RelReznikov}) for the classes of
the morphism $u$.

%-------------------------------------------------------------------------------------

\section{Preliminaries}\label{prelim}
%-------------------------------------------------------------------------------------

We briefly recall the definitions and terminologies from \cite{Le} 
and from the notes by Deligne
and Morgan \cite{De-Mo}.

Let $\cC^\infty$ be the sheaf of $C^\infty$-functions on $\R^p$. The space $\R^{p|q}$ is the topological space $\R^p$, endowed with the sheaf $\cC^\infty[\theta_1,...,\theta_q]$ of supercommutative super $\R$-algebras, freely generated over $\cC^\infty$ by the anticommuting $\theta_1,...,\theta_q$.
The coordinates $t^i$ of $\R^p$ and the $\theta_j$ and all generators of $\cC^\infty$ obtained from them by any automorphism are said to be the coordinates of $\R^{p|q}$. A supermanifold $M$ of dimension $p|q$ is a topological space $M_B$ (or also called as the body manifold with the structure sheaf $\cC^\infty_M$) endowed with a sheaf of super $\R$-algebras which is locally isomorphic to $\R^{p|q}$. The structure sheaf of $M$ is denoted by $\cO_M$.
We denote $p|q$, the real dimension of the supermanifold $M$.

On $M=\R^{p|q}$, the even derivations $\partial/\partial t^i$ and the odd derivations $\partial/\partial\theta^j$ are defined.

\begin{proposition}\cite[2.2.3]{Le}
The $\cO_M$-module of $\R$-linear derivations of $\cO_M$ is free of dimension $p|q$, with basis: the
 $\partial/\partial t^i$ and the $\partial/\partial\theta^j$.
\end{proposition}

Complex supermanifolds are topological spaces endowed with a sheaf of super $\comx$-algebras, locally isomorphic to some $(\comx^p,\cO[\theta^1,...,\theta^q])$. Here $\cO$ is the sheaf of holomorphic
functions on $\comx^p$. As before we denote $p|q$, the complex dimension of the complex supermanifold $M$.

Suppose $R$ be a commutative superalgebra and the standard free module $A^{r|s}$ is the module freely generated by even elements $e_1,...,e_r$ and odd elements $f_1,...,f_s$. 
An automorphism of $A^{r|s}$ is represented by an invertible matrix 
\begin{eqnarray}\label{matrix}
X = \left(  
\begin{array}{cc}
X_1 & X_2 \\
X_3 & X_4
\end{array}
\right) 
\end{eqnarray}
such that the $(r\times r)$--matrix $X_1$ and the $(s\times s)$--matrix $X_4$ have even entries and the $(s\times r)$--matrix $X_3$ and the $(r\times s)$--matrix $X_2$ have odd entries.
The group of all automorphisms of $A^{r|s}$ is denoted by $GL(r,s)$.

Suppose $M$ is a supermanifold and locally it looks like $\R^{p|q}$ as above. 
A \textit{complex supervector bundle} $V$ on $M$ is a fiber bundle $V$ over $M$ with typical fiber $\comx^{r|s}$ and structural group $GL(r,s)$. Alternately, it can be considered as a sheaf of $\cO_M$-supermodules  $\cV$, locally free of rank $r|s$. 

The tangent bundle $\cT_M$ is the $\cO_M$-module of derivations of $\cO_M$ and is a supervector bundle of rank $p|q$. The cotangent bundle $\Omega^1_M$ is the dual of $\cT_M$.
There is a differential $d:\cO_M\lrar \Omega^1_M$, giving rise to the super de Rham complex $\Omega^\bullet_M$ on $M$.

\begin{lemma}{\rm{(Poincar\'e lemma)}}\cite[p.73]{De-Mo}
The complex $\Omega^\bullet_M$ is a resolution of the constant sheaf on the body manifold $M_B$.
\end{lemma} 

In particular, the cohomology of $M_B$ can be computed by the super de Rham complex:
$$
H^*(M_B,\R)=H^*(\Gamma(M,\Omega^\bullet_M)).
$$

We briefly review the group of differential characters and Chern-Simons classes on a smooth manifold $X$.

\subsection{Analytic differential characters on $X$ \cite{Ch-Si}}\label{diff.char}

Let $S_k(X)$ denote the group of $k$-dimensional smooth singular chains on $X$, with integer coefficients. Let $Z_k(X)$ denote the subgroup of cycles.  Let us denote 
$$
S^\bullet(X,\Z):=\m{Hom}_\Z(S_\bullet(X),\Z)
$$
the complex of $\Z$ -valued smooth singular cochains, whose boundary operator is denoted by $\delta$.
The group of smooth differential $k$-forms on $X$ with complex coefficients is denoted by $A^k(X)$
and the subgroup of closed forms by $A^k_{cl}(X)$.
Then $A^\bullet(X)$ is canonically embedded in $S^\bullet(X)$, by integrating forms against the smooth
singular chains. In fact, we have an embedding
$$
i_\Z:A^\bullet(X)\hookrightarrow S^\bullet(X,\comx/\Z).
$$

The group of differential characters of degree $k$ is defined as
$$
\widehat{H^{k}}_\comx(X):=\{(f,\al)\in \m{Hom}_\Z(Z_{k-1}(X),\comx/\Z) \oplus A^k(X): \delta(f)=i_\Z(\al) \m{ and } d\al=0 \}.
$$
There is a canonical and functorial exact sequence:
\begin{equation}\label{diffeqn}
0\lrar H^{k-1}(X,\comx/\Z)\lrar \widehat{H^{k}}_\comx(X)\lrar A^k_{\Z}(X)\lrar 0.
\end{equation}
 
Here $A^k_{\Z}(X):=\m{ker}(A^k_{cl}(X)\lrar H^k(X,\comx/\Z))$.

Similarly, one can define the group of differential characters $\what{H^{k}}_\R(X)$ 
which have $\R/\Z$-coefficients.

%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%

\subsection{Cheeger-Chern-Simons classes}\label{Secondary classes}

Suppose $(E,\theta)$ is a vector bundle with a connection on $X$.
Then the characteristic forms
$$
c_k(E,\theta) \in A^{2k}_{cl}(X,\Z)
$$
for $0\leq k\leq r$=rank $(E)$, are defined using the universal Weil homomorphism \cite{Chn-Si}.
 
The characteristic classes 
$$
\widehat{c_k}(E,\theta) \in \widehat{H^{2k}}_\comx(X)
$$
are defined in \cite{Ch-Si} using a factorization of the universal Weil homomorphism. These classes are functorial lifting of the forms $c_k(E,\theta)$.

Similarly, there are classes
$$
\widehat{c_k}(E,\theta) \in \widehat{H^{2k}}_\R(X).
$$

\begin{remark}\label{lift}
If the forms $c_k(E,\theta)$ are zero, then the classes $\widehat{c_k}(E,\theta)$ lie in the cohomology $ H^{2k-1}(X,\comx/\Z)$. If $(E,\theta)$ is a flat bundle, then $c_k(E,\theta)=0$ and the classes $\widehat{c_k}(E,\theta)$ are called as the Chern-Simons classes of $(E,\theta)$. Notice that the class depends on the choice of $\theta$.
\end{remark}

Beilinson's theory of universal Chern-Simons classes yield classes for a flat connection $(E,\theta)$,
$$
\widehat{c_k}(E,\theta) \in H^{2k-1}(X,\comx/\Z)
$$
which are functorial and additive over exact sequences (see \cite{DHZ} and \cite{Esnault} for another construction).

%----------------------------------------------------------------------------------------

\section{Chern-Simons classes of flat connections on supermanifolds}

%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%

Let $(M,\cO_M)$ be a complex supermanifold of dimension $p|q$.
Consider the sheaf of differentials $\Omega^1_M$ 
on $M$ and let $\cE^{r|s}$ be a complex supervector bundle on $M$ of rank $r|s$.

\begin{lemma}\label{vectorbundle}
Given a complex supervector bundle $\cE^{r|s}$ of rank $r|s$ on $M$, there exists 
a direct sum decomposition
$$
E=E_0\oplus E_1
$$
for some complex smooth vector bundles $E_0$ and $E_1$ of rank $r$ and rank $s$ respectively, on the underlying $\cC^\infty$-manifold $M_B$.
\end{lemma} 
\begin{proof}
A rank $r|s$ complex supervector bundle $\cE^{r|s}$ determines two complex smooth bundles $E_0$ and $E_1$ on the underlying smooth manifold $M_B$ of $M$ as follows. One considers the body map
$$
\cO_M\lrar \cC^\infty_M\otimes \comx
$$
which is obtained by forgetting the local anticommuting variables $\theta_j$. Let $\ov{\cE^{r|s}}$ denote the sheaf of (super)sections of $\cE^{r|s}$. Then $\ov{E_{r+s}}:=\ov{\cE^{r|s}}\otimes_{\cO_M} (\cC^\infty_M\otimes \comx)$ is the sheaf of sections of a rank $r+s$ smooth 
complex vector bundle $E_{r+s}$ on the body manifold $M_B$.
Locally, the sheaf $\cE^{r|s}$ is generated by $r$ even elements and $s$ odd elements as a 
$\cO_M=\cC^\infty_M[\theta_1,...,\theta_q]$--module. Hence tensoring with $\cC^\infty_M$ locally
gives a rank $r+s$ free $\cC^\infty_M$--module given by the same generators.
This implies that the complex vector bundle $E_{r+r}$ is of rank $r+s$.
Now, we notice that the structural group of $\cE^{r|s}$ is $GL(r,s)$ and the structural group of the vector bundle $E_{r+s}$ factors via the projection
$$
GL(r,s)\rar GL(r+s).
$$
Using the description of the elements in $GL(r,s)$ in \eqref{matrix}, it follows that
the image under this projection consists of block diagonal matrices of sizes $r\times r$ and $s\times s$. 

This implies that the matrix of the transition functions of $E_{r+s}$ is a block diagonal 
matrix of rank $r$ and rank $s$ which correspond to smooth complex vector bundles $E_0$ and $E_1$ such that $r=\m{rank }E_0$ and $s=\m{rank }E_1$ on $M_B$.
\end{proof}

In view of the above lemma, we may regard a supervector bundle $\cE^{r|s}$ on $M$ as a supervector bundle
$E=E_0\oplus E_1$, on the underlying $\cC^\infty$-manifold $M_B$ where $E_0$ and $E_1$ are $\cC^\infty$-vector bundles on $M_B$.
 
\subsection{Superconnections}
Let $\cE^{r|s}= \,E=\,E_0\oplus E_1$ be a complex supervector bundle on a manifold $M_B$. Let $\Omega(M_B)=\oplus \Omega^p(M_B)$ be the algebra of smooth differential forms with complex coefficients.
Let
$$
\Omega(M_B,E):=\Omega(M_B) \otimes_{\Omega^0(M_B)}\Omega^0(M_B,E).
$$
where $\Omega^0(M_B,E)$ is the space of (super)sections of $E$.

Then $\Omega(M_B,E)$ has a grading with respect to $\Z\times \Z_2$

A \textit{superconnection} $D$ on $\cE^{r|s}$ is an operator on $\Omega(M_B,E)$ of odd degree
satisfying the derivation property
$$
D(\omega\alpha)=(d\omega)\alpha +(-1)^{deg\,\omega}\nabla\alpha.
$$
For example, a connection on $E$ preserving the grading when extended to an operator on $\Omega(M_B,E)$ in the usual way determines a superconnection.

The \textit{curvature} of a superconnection $D$ is the even degree operator $D^2:= D\circ D$ on $\Omega(M_B,E)$.

A superconnection is said to be \textit{flat} if $D^2=0$. We call the pair $(\cE^{r|s},D)$ as a flat complex supervector bundle.

We want to define Chern-Simons classes of $(E,D)$ when $D$ is a flat superconnection.
It is not immediately clear that $D$ induces a flat connection on the individual bundle $E_0$ and $E_1$.

For this purpose, we look at the situation, considered by Quillen \cite{Quillen}.

%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
\subsection{Quillen's construction}

Suppose $M$ is a supermanifold and $E$ is a complex supervector bundle on $M$. Regard $E=E_0\oplus E_1$ as a complex supervector 
bundle on $M$ in view of Lemma \ref{vectorbundle}, where $E_0$ and $E_1$ are smooth vector bundles on the body manifold $M_B$. Under this identification we omit the suffix $B$ from $M_B$ and without confusion we write $M=M_B$ in the following discussion.

Suppose $E$ is equipped with a superconnection $D$.
From the curvature $D^2$, one can construct differential forms
$$
\m{str}(D^2)^n= \m{str}D^{2n}
$$
in $\Omega(M)^{even}$. Here $\m{str}$ denotes the supertrace.
These are even forms since the supertrace preserves the grading. 

We have,
\begin{theorem}\label{quillen}
The form $\rm{str}D^{2n}$ is closed, and its de Rham cohomology class is independent of the choice of superconnection $D$.
\end{theorem}
\begin{proof}
See \cite[Theorem, p.91]{Quillen}.
\end{proof}

Quillen described the (super) Chern character of $E$ in terms of the usual Chern characters of $E_0$ and $E_1$ in the following situation.

Regard $E=E_0\oplus E_1$ as a complex supervector bundle and $D_0=D^0+D^1$ be a connection on $E$ preserving the grading. Let $L$ be an odd degree endomorphism of $E$ and write $D_t:=D+t.L$ where $t$
is a parameter.

\begin{proposition}$\rm{(Quillen)}$\cite{Quillen}\label{Quillen}
Replacing $L$ by $t.L$, where $t$ is a parameter, one obtains a family of forms 
\begin{equation}\label{traceform}
\rm{str}\,e^{(D+tL)^2}\,=\,\rm{str}\,e^{r^2L^2+ t[D,L]+ D^2}
\end{equation}
all of which represent the Chern character $\rm{ch}(E_0)-\rm{ch}(E_1)$ in the de Rham cohomology of $M$.

\end{proposition}
\eop

The referee has pointed out that the above computations on a supermanifold produces pseudodifferential forms. For our purpose, it suffices to note that the trace form in \eqref{traceform} defines a closed differential form whose de Rham class is independent of the superconnection \cite[Theorem p.91]{Quillen}.

 In this situation we want to define uniquely determined
 Chern-Simons classes of $(E,D_t)$ which is independent of $t$ and if $D^0, D^1$ are flat connections.
For this purpose, we look at the \textit{Character diagram} of Simons and Sullivan (see \cite{Sm-Su}):

\begin{equation}\label{Characterdiagram}
\begin{array}{ccccccc}
0 & & & & & & 0 \\
 &&&&&& \\
  & \searrow & & & & \nearrow &  \\
 &&&&&& \\
  &          & H^{k-1}(\R/\Z) & \sta{-B}{\lrar} & H^k(\Z) & & \\
&&&&&& \\
\searrow & \,\,\,\,\,\, \alpha{\nearrow} &\,\,\,\,\,\,\,\, \searrow i_1 & \,\,\,\,\,\,\,\,\,\,\,\delta_2\nearrow &\,\,\,\,\,\,\,\,\,\, \searrow r  & \,\,\,\,\,\,\,\,\,\nearrow \\
&&&&&& \\
& H^{k-1}(\R) & &\what{H^k}_\R(M)  && H^k(\R) & \\
&&&&&& \\
\nearrow &\,\,\,\,\,\,\,\,\, \searrow \beta &\,\,\,\,\,\,\,\,i_2 \nearrow &\,\,\,\,\,\,\,\,\,\,\,\,\, \searrow \delta_1 &\,\,\,\,\,\,\,\,\,\,\,\,\,\,s \nearrow &\,\,\,\,\,\,\,\,\,\,\,\,\, \searrow \\
&&&&&&\\
         &          & A^{k-1}/A^{k-1}_{\Z} & \sta{d}{\lrar}& A^k_{\Z} && \\
&&&&&& \\
 & \nearrow &&&& \searrow &   \\
&&&&&&\\
0 &&&&&& 0
\end{array}
\end{equation}

The diagonal sequences are exact and $(\alpha,B,r)$ is the Bockstein long exact sequence associated to the coefficient sequence $\Z\rar \R\rar \R/\Z$. Also $(\beta,d,s)$ is another long exact sequence in which $\beta$ and $s$ are defined via the de Rham theorem.
(A similar diagram holds by replacing $\R$ with $\comx$).

\begin{lemma}\label{unique}
Suppose $(F,\nabla)$ is a smooth flat connection on a manifold $M$. Then the Chern class
$c_i(F)\in H^{2i}(M,\Z)$ vanishes in $H^{2i}(M,\R)$. There is a unique lifting 
$\what{c_i}(F,\nabla)\in H^{2i-1}(M,\R/\Z)$ of the  integral class $c_i(F)$, for $i>0$.
\end{lemma}
\begin{proof}
The first assertion is by the Chern-Weil theory. The second assertion is by the Chern-Simons-Cheeger construction \cite{Ch-Si}.
\end{proof}

Consider the total Chern class
$$
c(F)= 1+c_1(F)+...+c_f(F)
$$
and the total Segre class
$$
s(F)=1+s_1(F)+...+s_f(F).
$$
Then we have the relations
\begin{equation}\label{segre}
s(F)=\f{1}{c(F)}
\end{equation}
and 
\begin{equation}\label{sum}
c(F-G)=c(F).s(G)
\end{equation}
where $G$ is any other vector bundle.
These relations also hold if we replace the classes $c_i(F)$ by $\what{c_i}(F,\nabla)$ and $s_i(F)$ by $\what{s_i}(F,\nabla)$ which are defined by the relation \eqref{segre}.
See \cite[p.64-65]{Ch-Si}.

To define a canonical lifting in $\what{H^*_\R}(M)$ of the trace form \eqref{traceform} associated to the superconnection $(E,D_t)$, one would need a notion of universal superconnection analogous to the universal connections defined by Narasimhan and Ramanan \cite{Narasimhan}, \cite{Narasimhan2}, which we may look in the future.

Nonetheless, we consider the family of superconnection $D_t=D+t.L$ as above. Consider the product manifold $\R\times M$ and the pullback $pr_2^*E$ of the bundle $E$. This bundle is equipped with a superconnection
$$
\bar{D}:= dt\partial_t+D'
$$
whose restriction to $\{t\}\times M$ is $D_t$. In terms of local trivialization of $E=M\times V$ we can describe
$\bar{D}$, $D'$ as follows. Write $D_t=d_M+\theta_t$, where $\theta_t$ is a family of one-forms on $M$ with values in $\m{End} V$ and let $\theta$ be the form on $\R\times M$ not involving $dt$ and having the restriction $\theta_t$ on $\{t\}\times M$. Then $D'=d_M+\theta$ and 
$$
\bar{D}=dt\partial_t+D'=(dt\partial_t+d_M)+\theta=d_{\R\times M} +\theta.
$$
See also \cite[p.91]{Quillen}.

By the homotopy property of de Rham cohomology, it follows that the class of $\m{str}D^{2n}_t$ in $H^{2n}(M,\R)$ is independent of $t$. 

\begin{proposition}\label{ChernSimons}
Suppose the superconnection $D_0=D^0\oplus D^1$ on the supervector bundle $E=E_0\oplus E_1$ preserves the grading and the individual smooth connections $D^0$ and $D^1$ are flat connections on $E_0$ and $E_1$ respectively. Then $D_t=D_0+t.L$ is a superconnection on $E$ where $L$ is an odd endomorphism of $E$. 
Then there is a uniquely determined class $\what{c_n}(E,D_t) \in H^{2n-1}(M,\R/\Z)$, independent of $t$. Moreover this class is equal to 
$$
\what{c_n}(E,D_0+L)=\sum_{p+q=n}\what{c_p}(E_0,D^0).\what{s_q}(E_1,D^1)
$$
\end{proposition}  
\begin{proof}
We notice that the trace forms are integral valued. This implies that the Chern class associated
to these forms lies in the integral cohomology $H^{2n}(M,\Z)$ which is independent of $t$ in
$H^{2n}(M,\R)$, by Quillen's Theorem \ref{quillen}. This determines a class in $H^{2n}(M,\Z)$ independent of $t$. But this class vanishes in $H^{2n}(M,\R)$ since $D_0$ has components $D^0$ and $D^1$ which are flat connections, hence $D_0^2=0$. Using the Bockstein operator in \eqref{Characterdiagram}, we conclude that there is a class
$$
\what{c_n}(E,D_t) \,\in\,H^{2n-1}(M,\R/\Z)
$$
which is independent of $t$ and we denote this class by $\what{c_n}(E,D_0+L)$.
We get a uniquely determined class $\what{c_n}(E,D_0+L)=\what{c_n}(E, D_0)\in H^{2n-1}(M,\R/\Z)$, as follows.

To get an expression for this class, we notice that by Quillen's result Proposition \ref{Quillen} the (super) Chern character form of $E$ is the difference $\m{ch}(E_0)-\m{ch}(E_1)$ in the integral cohomology. In particular we want to lift the integral Chern classes of $E_0-E_1$ in the $\R/\Z$-cohomology.
The relations in \eqref{segre} and \eqref{sum} give the formula
$$
\what{c_n}(E,D_0+L):=\sum_{p+q=n}\what{c_p}(E_0,D^0).\what{s_q}(E_1,D^1).
$$
The uniqueness of $\what{c_n}(E,D_0+L)$ follows from the uniqueness of the 
Chern-Simons classes $\what{c_p}(E_0,D^0)$ and $\what{c_q}(E_1,D^1)$, see Lemma \ref{unique}.
This concludes the lemma.
\end{proof}

\begin{remark}\label{sclasses}
All the above constructions follow verbatim by replacing $\R/\Z$-coefficients with $\comx/\Z$-coefficients. We call the resulting classes $\what{c_n}(E,D_0+L)\in H^{2n-1}(M,\comx/\Z)$ as the Chern-Simons classes of the superconnection $D_0+L$ (or $D_0+t.L$ for a parameter $t$).
\end{remark}

%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
\subsection{Chern Simons classes for a morphism between flat connections}

 Consider a homomorphism $u:E_0\rar E_1$ between complex vector bundles on a smooth manifold $M$. Then $u$ determines a class in the $K$-group $K(M)$.

Let
\begin{eqnarray}\label{L}
L=i \left(  
\begin{array}{cc}
0 & u^* \\
u & 0
\end{array}
\right) .
\end{eqnarray}
Here $u^*$ is the adjoint of $u$ relative to a given metric (see \cite{Quillen}). 
Regard $E=E_0\oplus E_1$ as a complex supervector bundle on $M$ and $D_0=D^0+D^1$ be a superconnection on $E$ preserving the grading. Then $L$ is an odd degree endomorphism of $E$ and as shown in \cite{Quillen}, $D+L$ is a superconnection and its curvature form $F=(D+L)^2$ is an even form with values in $\m{End}E$.

\begin{lemma}\label{relative}
Suppose $(E_0,D^0)$ and $(E_1,D^1)$ are flat connections and $u$ and $L$ are as above. 
Then we can define the Chern-Simons classes of the morphism $u$ (which need not be a flat morphism) 
in the $\comx/\Z$-cohomology of $M$ by setting
\begin{equation}\label{CSmorphism}
\what{c_n}(u)\,:=\,\what{c_n}(E,D_0+L)  \in\,H^{2n-1}(M,\comx/\Z)
\end{equation}
for $n\geq 1$, where $\what{c_n}(E,D_0+L)$ are defined in Lemma \ref{ChernSimons}.
\end{lemma}
\begin{proof}
The assumptions of Proposition \ref{ChernSimons} are fulfilled and we obtain uniquely defined classes
$$
\what{c_n}(u)\,:=\, \what{c_n}(E,D_0+L)  \in\,H^{2n-1}(M,\comx/\Z)
$$
for $n\geq 1$.
\end{proof}

\begin{theorem}$\rm{(Relative \,Reznikov\,theorem)}$\label{RelReznikov}
Suppose $u:E_0\rar E_1$ is a (not necessarily flat) homomorphism between flat complex vector bundles $(E_0,D^0)$ and $(E_1,D^1)$ on a compact K\"ahler manifold $M$ or a smooth complex quasi--projective variety $M$. Then the classes 
$$
\what{c_i}(u) \in H^{2i-1}(M,\comx/\Q)
$$ 
are zero, for $i\geq 2$. 
\end{theorem}
\begin{proof}
By Proposition \ref{ChernSimons}, Remark \ref{sclasses} and Lemma \ref{relative} we have the explicit expression of the class 
$$
\what{c_n}(u)=\sum_{p+q=n} \what{c_p}(E_0,D^0).\what{s_q}(E_1,D^1).
$$
When $M$ is a compact K\"ahler manifold then Reznikov's theorem \cite{Reznikov2} says that
$$
\what{c_n}(E_0,D^0),\, \what{c_n}(E_1,D^1) \in H^{2n-1}(M,\comx/\Z)
$$
are torsion, if $n\geq 2$. A similar result is true if $M$ is a smooth complex quasi--projective variety, by \cite{Iy-Si}. Since the classes $\what{s_q}$ are expressed in terms of $\what{c_i}$ for $i\leq q$, the assertion follows.
This proves the theorem
\end{proof}

{\Small
\textbf{Acknowledgements}:
We thank D. Leites for his comments and pointing out useful 
references to us, especially \cite{Quillen}, \cite{Quillen2}. His comments prompted us to improve
the results and understanding of this subject.
 The first named author is supported by the 
National Science Foundation (NSF) under agreement No. DMS-0111298.  
}

%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%

\end{document}